OpenStudy (anonymous):

Match each percent problem with its solution. WILL FAN AND MEDAL 1. 8 is 4% of what number? 2. 4 is what percent of 8? 3. What is 32% of 800 4. 8% of what number is 32? a. 50 b. 400 c. 256 d. 200

OpenStudy (anonymous):

i will fan and medal

OpenStudy (anonymous):

are you still here @miamoore

jhonyy9 (jhonyy9):

8 ----- 4% x ----- 100% ------------- x = 800/4 = 200 x=200

jhonyy9 (jhonyy9):

2. 4/8 = 1/2 so 4 is 50% of 8

jhonyy9 (jhonyy9):

3. 800 ------100% x -------- 32% ---------------- x = 800*32/100 = 8*32 =256 x = 256

jhonyy9 (jhonyy9):

4. 32 ----- 8% x ------ 100% --------------- x = 32*100/8 = 400 x=400

OpenStudy (anonymous):

thanks @jhoneyy9